two balls are drawn in succession without replacement from a jar containing 7 red balls and 6 blue balls. Let B be the random variable representing the number of blue balls. Find the mean.


1
Expert's answer
2022-03-24T03:40:37-0400

Cases

P(R,R)=7/(6+7) x 6/(6+6)=0.54x0.5=0.27

P(B,B)=6/(6+7)x5/(5+7)=0.46x0.42=0.19

P(R,B)=7/(6+7)x6/(6+6)=0.27

P(B,R)=6/(6+7)x7/(6+6)=0.27

"\mu=0\times0.27+2\times0.19+1\times0.27+1\times0.27=0.92"